OpenStudy (anonymous):

Show all work. 1) The sum of two integers is at most -10. One integer is 23. What is the other integer? 2) The difference between two integers is at least 12. The smaller integer is 2. What is the larger integer? 3) The sum of integer a and b is 13. The sum of integer a and integer c is 20. If integer b is 4, what is integer c? 4) Sherry bought a bag? It was on sale for $25. She saved $6. What was the original price? 5) Harry’s score on his test was 83. This was 7 points more than his friend’s

OpenStudy (anonymous):

1. x+23<=-10 x <= -33 2. x-2 >= 12 x>= 14 3.a+4 = 13 = a = 9 a+c = 20 9 + C = 20 c = 11 4. X - 6 = 25 X = 31 5. X + 7 = 83 x = 76